OpenStudy (anonymous):

P(600 less than or equal to X less than or equal to 640) mean=630 standard deviation= 7.94 --> p(a

OpenStudy (anonymous):

\[\begin{align*}P\left(600\le X\le640\right)&=P\left(\frac{600-630}{7.94}\le\frac{X-630}{7.94}\le\frac{640-630}{7.94}\right)\\\\ &=P\left(-3.778\le Z\le1.259\right) \end{align*}\]

OpenStudy (anonymous):

-3.778 is not on the chart though thats my problem

OpenStudy (anonymous):

Right, the probability is really, really small, so it's essentially the same as \(P(Z\le1.259)\).

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw